# Gyrokinetic flux-driven simulations in mixed TEM/ITG regime using a delta-f PIC scheme with evolving background

### v1

@@ V.2 Evolved profiles comparison
We now consider evolved f.s.a. profiles time averaged in the first time window $c_{s}t/a\in[592,613]$. All profiles are derived from binning marker $f_{0}\Omega$ values and $w$-weights into radial bins for the contributions from the background $f_{0}$ and perturbed $\delta f$ distributions respectively. Figure 9 shows the gyrocenter temperature logarithmic gradient for all cases considered. We first note the dip to zero in logarithmic gradients for both ions and electrons at $s=0.35$, which is expected to translate to the magnetic axis at later times as the heat is transported from the heat source peak towards the core, thus flattening the temperature profiles in this region (also see Fig.10). Figure 9(a) shows a steepening of gradient in $s\in[0.9,1.0]$ (well within the buffer region) for the ions, which is not present for the electrons. On the other hand, Fig. 9(b) shows a corrugation of the electron logarithmic gradient in the vicinity of the $q=1$ flux surface at $s\approx 0.55$. It can be further seen in the inserted figure in Fig. 9(b), that in the adaptive cases, the background electron temperature also exhibits such corrugation, illustrating the ability of the evolving background to capture such fine profile features. No such corrugation is observed for the ions. Such corrugated structures are related to those of the $\omega_{E\times B}$ radial profiles of Fig. 8. Finally, we see that for both species, the adaptive cases resulted in a lower logarithmic gradient in region $s\in[0.6,0.9]$. Despite this fact, adaptive cases still exhibit greater local heat diffusivity values (see Fig.3(a)) due to lower local $\omega_{E\times B}$ amplitudes, as mentioned.
(a) Ion
(b) Electron
@@
Figure 9: Time averaged $c_{s}t/a\in[592,613]$ f.s.a. gyrocenter temperature logarithmic gradient profiles for various marker numbers $N_{p}$ under the standard and adaptive schemes. Dashed and solid lines show contribution by $f_{0}$ and $f_{0}+\delta f$ respectively.
@@
Ion and electron temperature profiles are shown in Fig.10. We first note that the temperature in the core has risen more under the standard scheme as compared to the adaptive scheme. All profiles converged in marker number $N_{p}$ under the respective schemes. Note that while at this point in time, there is no off-axis peak ion temperature profile, there is one for the electron temperature under the standard scheme, at around $s=0.45$. This does not coincide with the peak of the heat source for the electrons at around $s=0.55$ (see Fig.2(b)). The slight corrugation for the electron temperature at $s=0.55$ was reflected in its logarithmic gradient of Fig. 9(b). Next, there is an increase in ion temperature at the magnetic axis under the standard scheme, with the $N_{p}=128$M case having a larger increase than that of the $N_{p}=256$M. A slight increase in ion temperature at the magnetic axis under the adaptive scheme is also observed. We thus attribute this problem as numerical, and is related to the under-sampling of phase space volume. This is suggested by the dip in background temperature profiles (dashed) of the standard cases (black and orange). As these background profiles are not adaptive, we expect them to coincide with the initial profile (dashed gray). Between the standard cases, the case with $N_{p}=128$M deviates more from its initial values compared to the case with $N_{p}=256$M. This effect is not observed for electrons. This sampling problem of phase-space volume will be discussed in Sec.V.3.
(a) Ion
(b) Electron
@@
Figure 10: Time averaged $c_{s}t/a\in[592,613]$ f.s.a. gyrocenter temperature profiles for various marker numbers $N_{p}$ under the standard and adaptive schemes. Dashed and solid lines show contribution by $f_{0}$ and $f_{0}+\delta f$ respectively.
@@
Figure 11 shows the time-averaged gyrocenter densities of the ions and electrons. Focusing first on Fig. 11(a), we see that at the considered stage of the simulation, the non-adaptive background density of the standard cases (dashed black and orange) has a lower value near the magnetic axis as compared to the initial profile (dashed gray). The source of this deviation is the same as that of Fig. 10(a) for the ion temperature. Next, the $f_{0}$ contributed density of the adaptive cases (dashed blue and green) is on top of the one including the $\delta f$ contribution, which indicates that the time-dependent background density has captured the total f.s.a. density evolution correctly. Furthermore, the adaptive cases seem to have converged in marker number $N_{p}$, whereas the standard cases has not, with $N_{p}=256$M for the standard case showing a smaller deviation of the evolved density profile. Figure 11(b) shows the difference between ion gyrocenter and electron densities, where we see a larger difference between these densities for the standard cases. And between these standard cases, the case with $N_{p}=128$M shows a larger density difference as compared to that with $N_{p}=256$M, thus indicating that this density difference is partly due to sampling noise. The same $N_{p}$ trend can be observed for the adaptive cases, though the results are more converged. Ion gyrocenter and electron density difference is not expected to be zero, as one still needs to account for ion Finite-Larmor-Radius (FLR) effects and ion polarisation density. Taking a step back, we now see in Fig. 11(a) that the apparent convergence of the standard case with $N_{p}=128$M (solid orange) with the adaptive cases (solid blue and green) is just a coincidence.
@@
Figure 11: Time averaged $c_{s}t/a\in[592,613]$ (a) gyrocenter density profiles and (b) ion-electron gyrocenter density difference, for various marker numbers $N_{p}$ under the standard and adaptive schemes. Dashed and solid lines show contribution by $f_{0}$ and $f_{0}+\delta f$ respectively.
@@ V.3 Phase-space volume sampling
Figure 12: Ion phase-space volume diagnostic applied to Cartesian bins in $(s,v^{2})$-space for the standard and adaptive cases for various marker numbers $N_{p}$ at $c_{s}t/a=612$. Shown is the ratio $\Omega(t)/\Omega^{\rm(a)}$, of which values near $1$ reflect good sampling, while ratios deviating significantly from $1$ reflect poor/deficient sampling. The red dashed line on the contour plots indicate the energy per mass upper-bound during initial marker loading, given by $\kappa_{v}^{2}v_{\rm th}^{2}(s)$. All plots share the same colour scale.
At the beginning of each simulation, markers are distributed uniformly in configuration space $\vec{R}=[s,{\theta^{\star}},\varphi]$ of volume enclosed by radius $s$
$V(s)$ $=$ $\int_{0}^{s}\,\text{d}^{3}R=\int_{0}^{s}\,\text{d}s\,\text{d}{\theta^{\star}}\,\text{d}\varphi\,J_{s}(s,{\theta^{\star}}),$ (19)
where $\,\text{d}^{3}R=J_{s}\,\text{d}s\,\text{d}{\theta^{\star}}\,\text{d}\varphi$ is the configuration space differential volume element, and prime denotes the derivative of its argument. In the velocity $(v_{\parallel},v_{\perp})$-space, markers are distributed uniformly in a semi-circular disk of radius $\kappa_{v}v_{\rm th}(s)$, $v_{\parallel}^{2}+v_{\perp}^{2}\leq\kappa_{v}^{2}v_{\rm th}^{2}(s)$, with $v_{\perp}\geq 0$. Here, $\kappa_{v}=5$ is set for all simulations of this work. As the simulation evolves over time, markers move in phase space governed by the equations of motion Eqs.(2), and the overall marker distribution should behave like an incompressible fluid. Due to time-integration inaccuracies, markers will deviate from their exact orbits, and phase space may not be well sampled with simulation time. This is particularly true near the magnetic axis $s=0$ where, according to the current loading, there are very few markers per $s$-bin. A useful diagnostic is to bin phase-space volume of each marker onto a grid of reduced $[s,v^{2}]$-phase-space. Let $(i,j)$ be the index of the bin $(s,v^{2})\in\Gamma_{ij}=[s_{i},s_{i+1}]\times[v_{j}^{2},v_{j+1}^{2}]$. Then, the loaded phase-space volume at initial time of the $(i,j)^{\rm th}$ bin is given by
@@
Where $V^{\prime}(s)$ is the derivative of the configuration space volume Eq.(19). Figure 12 shows the binned ion phase-space volume $\Omega_{ij}(t)=\sum_{p\in\Gamma ij}\Omega_{p}(t)$ normalized to Eq.(LABEL:eq:pvol) at time $c_{s}t/a=613$. A feature common to all cases is that there is a diffusion of sampled phase-space volume $\Omega$ around the velocity cut-off, as expected, with greatest diffusion around $s\in[0.5,0.6]$. Focusing now on the standard cases Figs.12(a)-12(b), we see an under-sampling in $\Omega$ around $s\in[0,0.3]$ at low energies. We also see that the situation is improved with higher $N_{p}$, indicating that this problem is of numerical origin. The same is true for the adaptive cases Figs.12(c)-12(d), though the extent of the under-sampling is much smaller compared to the standard cases. To determine if this phase-space under-sampling is due to marker migration or simply due to phase-space volume corruption, we consider now the marker distribution in $(s,v^{2})$-space. The expression for the loaded differential marker number $\,\text{d}N$ in differential area $\,\text{d}s\,\text{d}v^{2}$ at initial time is given by
@@
Figure 13: Ion marker count for different $s$-cuts for the standard and adaptive cases with $N_{p}=128$M at initial time (dashed) and at $c_{s}t/a=612$ (solid). Here, $N_{max}(t=0)$ is the global maximum marker per $(s,v^{2})$-bin count at initial time $t=0$.
Figure 13 shows the various radial cuts of the ion marker number per $(s,v^{2})$-bin for the plots Figs.12(b) and 12(c) respectively, at the same instance $c_{s}t/a=612$ and with $N_{p}=128$M. We note a general diffusion of markers towards lower energies. At the magnetic axis $s=0.02,0.10$ (black,red) however, there is greater marker diffusion to lower energies for the adaptive case compared to the standard case. Nonetheless, given that there are at least as many markers at $c_{s}t/a=612$ as at $t=0$, considering Fig.12(b), we conclude that the under-sampling near magnetic axis is the result of phase-space volume corruption. The smaller extent of the under-sampling in Fig.12(c) for the adaptive case seem to be compensated by the larger accumulation of markers Fig. 13(b). This however requires further investigation in a future work.
The problem of under-sampling for the standard cases Figs.12(a)-12(b) and 13(a) thus explains the density drop in Fig.11(a) and temperature rise in Fig.10(a). The former is due to the lack of markers being accounted for, and the latter is due to the lack of the contribution of low-energy markers. In terms of the gyrokinetic equations of motion Eqs.2, the difference between the standard and adaptive schemes lies in the evaluation of the $E\times B$ drift term involving the self-consistent gyroaveraged electrostatic potential $\phi$. The adaptive scheme allows for a more accurate evaluation of the r.h.s. of the QNE Eq.(LABEL:eq:qne_hyb) using a better control variate $f_{0}$, this leading to a more accurate solution for $\phi$. Moreover, under the standard scheme, a better evaluation of $\phi$ requires a higher $N_{p}$. Indeed, based on Figs.12(a)-12(b), the under-sampling situation at $s=0$ improves with higher $N_{p}$. However, the incompressibility of phase space is verified for any field $\phi$, be it self-consistent or not. Nonetheless, non-smooth solutions of $\phi$ as a result of poor integration can lead to spurious $E\times B$ drifts, which leads ultimately to the difference in marker trajectories between the standard and adaptive cases. Though not shown, no such under-sampling is observed for the electrons in both the standard and adaptive cases.
@@ V.4 Late time profiles
Figure 14: Time averaged $c_{s}t/a\in[1246,1266]$ (solid) and $c_{s}t/a\in[592,613]$ (dotted), f.s.a. gyrocenter temperature profiles for various marker numbers $N_{p}$ under the adaptive scheme. All colored curves show contribution by $f_{0}+\delta f$.
@@
This section describes profiles that have reached quasi-steady state with just one marker loading at initial time. For the given fluxes, only adaptive cases maintain adequate simulation quality SNR levels to reach this simulation time $c_{s}t/a=1266$ as shown in Fig.7. Figure 14 shows the temperature profiles for the ions and electrons separately, contributed by the respective total distribution $f_{0}+\delta f$. Profiles contributed only by the background $f_{0}$ are not shown as they coincide with the total contribution at this resolution, indicating that the adaptation rates are large enough and $\left\langle\delta f\right\rangle$ is very small. Also shown are the corresponding temperature profiles of the previous time window $c_{s}t/a=[592,613]$ (colored dashed). We see that f.s.a. temperature profiles at $c_{s}t/a=1266$ has risen mainly around $s=0.4$, resulting in a maximum deviation of about $20\%$ from its respective initial values. The increase in temperature at the magnetic axis for the ions, Fig.14(a), is due to the under-sampling problem discussed in Sec.V.3 and illustrated in Figs.12(c)-12(d). Except for the temperature rise at the magnetic axis, the profiles with $N_{p}=128$M and $N_{p}=64$M of the respective species seem to converge. The logarithmic gradients of these temperature profiles are shown in Fig.15. The risen peak in temperature around $s\in[0.3,0.4]$ has caused an increase and decrease of the $-\left\langle R/L_{T}\right\rangle$ values to the left and right of this region, respectively. Furthermore, the temperature peak has lead to negatives values of $-\left\langle R/L_{T}\right\rangle$. Both ions and electrons also have increased temperature gradients in the pedestal region $s\in[0.8,1.0]$. This is however not explained by the respective heat sources Fig.2 as this region is well within the buffer, albeit with a weak buffer strength $\gamma_{\rm b}=11\%\gamma_{\rm max}$. Logarithmic gradient values at the flat gradient region $s\in[0.6,0.9]$ on the other hand remains approximately constant in time after a decrease before the first time window $c_{s}t/a\in[592,613]$. The corrugation of the electron temperature logarithmic gradient profile around $s\approx 0.55$ has also remained relatively constant in amplitude.
@@
Figure 15: Time averaged $c_{s}t/a\in[1246,1266]$ (solid) and $c_{s}t/a\in[592,613]$ (dotted), f.s.a. gyrocenter temperature logarithmic gradient profiles for various marker numbers $N_{p}$ under the adaptive scheme. All colored curves show contribution by $f_{0}+\delta f$.
We now investigate the density profile evolution at quasi-steady state. Figure 16 shows the flux-surface- and time-averaged in $c_{s}t/a\in[1246,1266]$ (colored solid) density and its logarithmic gradient profiles for the ions and electrons contributed by $f_{0}+\delta f$ under the adaptive scheme. For comparison, the density values at the previous time window $c_{s}t/a\in[592,613]$ (colored dashed) are shown in Fig.16(a). As ion gyrocenter flux decreases approaching quasi-steady state, so does the change in ion gyrocenter density profiles, Fig.16(a). The adaptive cases with $N_{p}=128$M and $64$M seem to have converged density everywhere except near the magnetic axis, due to aforementioned ion under-sampling of phase-space. Though not shown, the difference between ion gyrocenter and electron density for the adaptive cases (blue and green) remains approximately the same between the two time-averaging windows $c_{s}t/a\in[592,613]$ and $c_{s}t/a\in[1246,1266]$, see Fig.11(b). Taken as a whole, species densities have deviated up to $50\%$ from their initial values. This amount of deviation would certainly challenge the delta-f PIC constraint of $\|\delta f\|/\|f\|\ll 1$. We consider now both ion gyrocenter (solid) and electron (dashed) density logarithmic gradient of the adaptive cases as shown in Fig.16(b). We first note the significant decrease of $50\%$ from their initial values in the radial region $s\in[0.4,1.0]$. For the region $s\in[0.2,0.4]$ however, there is a increase in $-\left\langle R/L_{n}\right\rangle$ values as the density profiles relax. We see once again the corrugation in the electron density logarithmic gradient near the $q=1$ flux surface which is absent for the ions, which has been observed in Figs.9(b) and 15(b) for electron temperature logarithmic gradient profiles.
@@
Figure 16: Time averaged $c_{s}t/a\in[1246,1266]$, f.s.a. (a) ion gyrocenter (solid) density and (b) ion gyrocenter (solid) and electron (dashed) density logarithmic gradient profiles, for various marker numbers $N_{p}$ under the adaptive scheme. Also shown in (a) is the ion gyrocenter density in the previous time window $c_{s}t/a\in[592,613]$ (dashed). All colored curves show contribution by $f_{0}+\delta f$.

### v3

@@ V.2 Evolved profiles comparison
(a) Ions
(b) Electrons
@@
Figure 9: Time averaged $c_{s}t/a\in[592,613]$ f.s.a. gyrocenter temperature logarithmic gradient profiles for various marker numbers $N_{p}$ under the standard and adaptive schemes for (a) ions and (b) electrons. Dashed and solid curves show contribution by $f_{0}$ and $f_{0}+\delta f$ respectively. Initial profiles given by dotted gray curve. Vertical dashed line indicates position of $q=1$ surface.
We now consider evolved f.s.a. profiles, further averaged over the first time window $c_{s}t/a\in[592,613]$ (shown e.g. in Fig. 3). All profiles are derived from binning marker $f_{0}\Omega$ values and $w$-weights into radial bins for the contributions from the background $f_{0}$ and perturbed $\delta f$ distributions respectively. Figure 9 shows the gyrocenter temperature logarithmic gradient for all cases considered and for both ions and electrons. We first note the dip towards zero in logarithmic gradients for both species inwards from the heat sources (see Fig. 2) at $s=0.35$. This dip moves further towards the magnetic axis at later times as the heat is transported from the heat source peak towards the core, thus flattening the temperature profiles in this region (also see Fig.10). Figure 9(a) shows a steepening of gradient in $s\in[0.9,1.0]$ (well within the buffer region) for the ions, which is barely present for the electrons (in Fig.9(b)). On the other hand, in the adaptive cases Fig. 9(b) shows a corrugation of the electron logarithmic gradient in the vicinity of the $q=1$ flux surface at $s\approx 0.55$. This can be further seen in the inserted figure in Fig. 9(b), showing a zoom-in of this region and illustrating the ability of the evolving background to capture such fine profile features. No such corrugation is observed for the ions. This corrugated structure is related to the one of the $\omega_{E\times B}$ radial profiles of Fig. 8. Finally, we see that for both species, the adaptive cases resulted in a lower logarithmic gradient in region $s\in[0.6,0.9]$. This is related to the fact that adaptive cases exhibit greater local heat diffusivity values (see Fig.3(a)) due to lower local $\omega_{E\times B}$ amplitudes, as mentioned.
@@
Ion and electron temperature profiles are shown in Fig.10. Note that while at this point in time, there is no off-axis peak ion temperature profile, there is one for the electron temperature under the standard scheme, at around $s=0.45$. This does not coincide with the peak of the heat source for the electrons at around $s=0.55$ (see Fig.2(b)). The slight corrugation for the electron temperature at $s=0.55$ was emphasized in its logarithmic gradient in Fig. 9(b). One further notes an increase in ion temperature at the magnetic axis under the standard scheme, with the $N_{p}=128$M case having a larger increase than that of the $N_{p}=256$M simulation. Only a slight increase in ion temperature at the magnetic axis under the adaptive scheme is observed. We thus attribute this problem as numerical, and is related to the under-sampling of phase space volume discussed in Sec. V.3. This is already suggested by the dip in background temperature profiles (dashed) of the standard cases (black and orange). As these background profiles are not adaptive, we expect them to coincide with the initial profile (dashed gray). Between the standard cases, the case with $N_{p}=128$M deviates more from its initial values compared to the case with $N_{p}=256$M. This effect is not observed for electrons. This sampling problem of phase-space volume will be discussed in Sec.V.3.
(a) Ions
(b) Electrons
@@
Figure 10: F.s.a. gyrocenter temperature profiles time averaged over $c_{s}t/a\in[592,613]$ for various marker numbers $N_{p}$ under the standard and adaptive schemes. Dashed and solid lines show contribution by $f_{0}$ and $f_{0}+\delta f$ respectively, essentially on top of each other in adaptive cases.
@@
Figure 11 shows the gyrocenter densities of the ions and electrons time-averaged over $c_{s}t/a\in[592,613]$. Focusing first on Fig. 11(a), we see that at the considered stage of the simulation, the non-adaptive background density of the standard cases (dashed black and orange) has a lower value near the magnetic axis as compared to the initial profile (dashed gray). The source of this deviation is the same as that in Fig. 10(a) for the ion temperature, i.e. under-sampling of phase space. Next, the $f_{0}$ contribution to density of the adaptive cases (dashed blue and green) is nearly identical to the one including the $\delta f$ contribution, which indicates that the time-dependent background density has correctly captured the evolution of the total f.s.a. density. Furthermore, the adaptive cases seem to have converged in marker number $N_{p}$, whereas the standard cases has not, with $N_{p}=256$M for the standard case showing a smaller deviation of the evolved density profile compared to the $N_{p}=128$M case. Figure 11(b) shows the difference between ion gyrocenter and electron densities, where we see a larger difference between these densities for the standard cases. Between these standard cases, the case with $N_{p}=128$M shows a larger density difference as compared to that with $N_{p}=256$M, thus indicating that this density difference is partly due to sampling error, as pointed out in Figs. 10 and 11(a). The same $N_{p}$ trend can be observed for the adaptive cases, though the results are clearly already more converged. Note that ion gyrocenter and electron density difference is not expected to be zero, as one still needs to account for ion Finite-Larmor-Radius (FLR) effects and ion polarisation density. Taking a step back, we now see in Fig. 11(a) that the apparent convergence of the standard case with $N_{p}=128$M (solid orange) with the adaptive cases (solid blue and green) is just a coincidence.
@@
Figure 11: (a) ion gyrocenter density profiles and (b) ion-electron gyrocenter density difference, time averaged over $c_{s}t/a\in[592,613]$ for various marker numbers $N_{p}$ under the standard and adaptive schemes. Dashed and solid lines show contribution by $f_{0}$ and $f_{0}+\delta f$ respectively.
@@ V.3 Phase-space volume sampling
Figure 12: Ion phase-space volume diagnostic applied to Cartesian bins in $(s,v^{2})$-space for the standard and adaptive cases for various marker numbers $N_{p}$ at $c_{s}t/a=612$. Shown is the ratio $\Omega(t)/\Omega^{\rm(a)}$. Values near $1$ reflect good sampling, while ratios deviating significantly from $1$ reflect poor/deficient sampling. The red dashed line on the contour plots indicates the energy per mass upper-bound during initial marker loading, given by $\kappa_{v}^{2}v_{\rm th}^{2}(s)$. All plots share the same colour scale.
The configuration space $\vec{R}=[s,{\theta^{\star}},\varphi]$ volume enclosed by the magnetic surface labeled by $s$ be given by
$V(s)$ $=$ $\int\,\text{d}^{3}R=\int_{0}^{s}\,\text{d}s\int_{0}^{2\pi}\,\text{d}{\theta^{\star}}\int_{0}^{2\pi}\,\text{d}\varphi\,J_{s}(s,{\theta^{\star}}),$ (19)
where $\,\text{d}^{3}R=J_{s}\,\text{d}s\,\text{d}{\theta^{\star}}\,\text{d}\varphi$ is the configuration space differential volume element, and prime denotes the derivative of its argument. At the beginning of each simulation, markers are distributed uniformly in $V(s=1)$. In the velocity $(v_{\parallel},v_{\perp})$-space, markers are distributed uniformly in a semi-circular disk of radius $\kappa_{v}v_{\rm th}(s)$, $v_{\parallel}^{2}+v_{\perp}^{2}\leq\kappa_{v}^{2}v_{\rm th}^{2}(s)$, with $v_{\perp}\geq 0$. Here, $\kappa_{v}=5$ is set for all simulations of this work. As the simulation evolves over time, markers move in phase space according to the equations of motion (2), and the overall marker distribution should behave like an incompressible fluid. Due to time-integration inaccuracies, markers will deviate from their exact orbit and as a result incompressibility is not exactly ensured, reflected by increasing errors in phase space sampling over time. This is particularly true near the magnetic axis $s=0$ where, according to the current loading, there are very few markers per $s$-bin. A useful diagnostic is to bin the phase-space volume $\Omega_{p}$ of each marker onto a grid of reduced $[s,v^{2}]$-phase-space. Let $(i,j)$ be the index of such a bin $\Gamma_{ij}=[s_{i},s_{i+1}]\times[v_{j}^{2},v_{j+1}^{2}]$. The exact phase space volume of the $(i,j)^{\rm th}$ bin is given by
@@
where $V^{\prime}(s)$ is the derivative of the configuration space volume Eq.(19). Figure 12 shows the binned ion phase-space volume $\Omega_{ij}(t)=\sum_{p\in\Gamma ij}\Omega_{p}(t)$ normalized to the reference value Eq.(LABEL:eq:pvol) at time $c_{s}t/a=613$. A feature common to all cases is that there is a diffusion of sampled phase-space volume $\Omega$ around the velocity cut-off at $\kappa_{v}v_{\rm th}$, as expected, with greatest diffusion around $s\in[0.5,0.6]$. Focusing on the standard cases, Figs.12(a)-12(b), we see an under-sampling in $\Omega$ around $s\in[0,0.3]$ at low energies. We also see that the situation is somewhat improved with higher $N_{p}$, indicating that this problem is of numerical origin. The same is true for the adaptive cases presented in Figs.12(c)-12(d), though the extent of the under-sampling is much smaller compared to the standard cases. To determine if this phase-space under-sampling is due to integration errors in marker trajectories or phase-space volume corruption, we consider now the marker distribution in $(s,v^{2})$-space. The expression for the loaded marker number $\,\text{d}N$ in the differential area $\,\text{d}s\,\text{d}v^{2}$ at initial time is given by
@@
(a) Standard
(b) Adaptive
Figure 13: Ion marker count for different $s$-cuts for the (a) standard and (b) adaptive cases with $N_{p}=128$M at initial time (dashed) and at $c_{s}t/a=612$ (solid). Here, $N_{max}(s,t=0)$ is the global maximum marker per $(s,v^{2})$-bin count at initial time $t=0$.
Figure 13 shows the various radial cuts of the ion marker number per $(s,v^{2})$-bin corresponding to Figs.12(b) and 12(c) respectively, i.e. at the same instant $c_{s}t/a=612$ and with $N_{p}=128$M. We note a general accumulation of markers towards lower energies. Near the magnetic axis at $s=0.02,0.10$ (black,red) however, there is greater marker accumulation to lower energies for the adaptive case compared to the standard case, thus concluding that the under-sampling near the magnetic axis is the result of phase-space volume corruption. The smaller extent of the under-sampling in Fig.12(c) for the adaptive case seems to be ensured by the larger accumulation of markers shown in Fig. 13(b). Further investigation in a future work will be required to fully understand this issue.
The problem of under-sampling for the standard cases Figs.12(a)-12(b) and 13(a) thus explains the density drop in Fig.11(a) and temperature rise in Fig.10(a). The former is due to the lack of markers being accounted for, and the latter is due to the lack of the contribution of low-energy markers. As temperature is the average energy over density, the under-estimated density also contributes to the over-estimation of temperature. In terms of the gyrokinetic equations of motion (2), the difference between the standard and adaptive schemes lies in the evaluation of the $E\times B$ drift term involving the self-consistent gyroaveraged electrostatic potential $\phi$. The adaptive scheme allows for a more accurate evaluation of the r.h.s. of the QNE (LABEL:eq:qne_hyb) using a better control variate $f_{0}$, this in turn leading to a more accurate solution for $\phi$. Moreover, under the standard scheme, a better evaluation of $\phi$ requires a higher $N_{p}$. Indeed, based on Figs.12(a)-12(b), the under-sampling issue near $s=0$ improves with higher $N_{p}$. Even though the incompressibility of phase space should be verified for any field $\phi$, be it self-consistent or not, non-smooth solutions of $\phi$ to (LABEL:eq:qne_hyb) as a result of poor (noisy) integration of the r.h.s. can lead to $E\times B$ drifts in the marker trajectories which are difficult to integrate in time thus violating phase space incompressibility. Though not shown, no such under-sampling is observed for the electrons in both the standard and adaptive cases.
@@ V.4 Late time profiles
Figure 14: F.s.a. gyrocenter temperature profiles time averaged over $c_{s}t/a\in[1246,1266]$ (solid) and $c_{s}t/a\in[592,613]$ (dotted) for various marker numbers $N_{p}$ under the adaptive scheme for (a) ions and (b) electrons. All colored curves show contributions by $f_{0}+\delta f$.
@@
This section describes profiles for simulations that have been able to reach quasi-steady state with just one marker loading at initial time. For the considered heat sources, only adaptive cases clearly maintain adequate SNR levels to reach the simulation time $c_{s}t/a=1266$ as shown in Fig.7. Figure 14 shows the temperature profiles for the ions and electrons separately, contributed by the respective total distributions $f_{0}+\delta f$. Profiles contributed only by the background $f_{0}$ are not shown as they closely follow the total contribution, indicating that the adaptation rates are large enough and as a result $\left\langle\delta f\right\rangle$ is very small. Also shown are the corresponding temperature profiles of the previous time window $c_{s}t/a=[592,613]$ (colored dashed). We see that f.s.a. temperature profile at $c_{s}t/a=1266$ has further risen mainly around $s=0.4$, resulting in this region in a maximum deviation of about $20\%$ from initial values. As already mentioned for Fig.10(a) the increase in temperature at the magnetic axis for the ions, Fig.14(a), is due to the under-sampling problem discussed in Sec.V.3 and illustrated in Figs.12(c)-12(d). Except for the temperature rise at the magnetic axis, the profiles with $N_{p}=128$M and $N_{p}=64$M of the respective species seem to be converged. The logarithmic gradients of these temperature profiles are shown in Fig.15. The risen peak in temperature around $s\in[0.3,0.4]$ has caused an increase and decrease of the $-\left\langle R/L_{T}\right\rangle$ values to the left and right of this region, respectively. Furthermore, the temperature peak has led to negative values of $-\left\langle R/L_{T}\right\rangle$. Both ions and electrons also have increased temperature gradients in the pedestal region $s\in[0.8,1.0]$. This is probably caused by edge buffer region and homogeneous Dirichlet condition at $s=1$ for $\phi$ both contribute to reduction of turbulence leading to profile steepening at the outer boundary. Logarithmic gradient values at the flat gradient region $s\in[0.6,0.9]$ on the other hand remains approximately constant in time after a decrease before the first time window $c_{s}t/a\in[592,613]$. The corrugation of the electron temperature logarithmic gradient profile around $s\approx 0.55$ has also remained relatively constant in amplitude.
(a) Ions
(b) Electrons
@@
Figure 15: F.s.a. gyrocenter temperature logarithmic gradient profiles time averaged over $c_{s}t/a\in[1246,1266]$ (solid) and $c_{s}t/a\in[592,613]$ (dotted), for various marker numbers $N_{p}$ under the adaptive scheme for (a) ions and (b) electrons. All colored curves show contributions by $f_{0}+\delta f$.
@@
Figure 16: F.s.a. (a) ion gyrocenter (solid) density and (b) ion gyrocenter (solid) and electron (dashed) density logarithmic gradient profiles time averaged over $c_{s}t/a\in[1246,1266]$, for various marker numbers $N_{p}$ under the adaptive scheme. Also shown in (a) is the ion gyrocenter density in the previous time window $c_{s}t/a\in[592,613]$ (dashed). All colored curves show contributions by $f_{0}+\delta f$.
We now investigate the density profile evolution at quasi-steady state. Figure 16 shows the flux-surface- and time-average over $c_{s}t/a\in[1246,1266]$ (colored solid) density and its logarithmic gradient profiles for the ions and electrons contributed by $f_{0}+\delta f$ under the adaptive scheme. For comparison, the density values at the previous time window $c_{s}t/a\in[592,613]$ (colored dashed) are shown in Fig.16(a). As the ion gyrocenter flux decreases approaching quasi-steady state, so does the change in ion gyrocenter density profiles given that the simulations include no particle sources, Fig.16(a). The adaptive cases with $N_{p}=128$M and $64$M seem to have converged density profiles everywhere except near the magnetic axis, due to aforementioned ion under-sampling of phase-space. The difference in the profiles between the two simulations with different $N_{p}$ is however very small ($5\%$).Though not shown, the difference between ion gyrocenter and electron density for the adaptive cases (blue and green) remains approximately the same between the two time-averaging windows $c_{s}t/a\in[592,613]$ and $c_{s}t/a\in[1246,1266]$, see Fig.11(b). Taken as a whole, species densities have decreased by $50\%$ from their initial values. This amount of deviation certainly challenges the delta-f PIC constraint of $\|\delta f\|/\|f\|\ll 1$ in the standard approach without background adaptation. We consider now both ion gyrocenter (solid) and electron (dashed) density logarithmic gradient of the adaptive cases as shown in Fig.16(b). We first note the significant decrease of $50\%$ from their initial values in the radial region $s\in[0.4,1.0]$. For the region $s\in[0.2,0.4]$ however, there is a local increase in $-\left\langle R/L_{n}\right\rangle$ as the actual density values decrease. We see once again the corrugation in the electron density logarithmic gradient near the $q=1$ flux surface which is absent for the ions, which has already been observed in Figs.9(b) and 15(b) for electron temperature logarithmic gradient profiles.